Nnesha (nnesha):

question!!!! :( You inherit $10,000 with the stipulation that for the first year the money must be invested in two stocks paying 6% and 11% annual interest, respectively. how much should be invested at each rate if the total interest earned for the year is to be $700? ^_^ :(

OpenStudy (misty1212):

HI!!

Nnesha (nnesha):

hai ^_~

Nnesha (nnesha):

i have answer method everything i need short and easy way to do this :(

OpenStudy (misty1212):

is this a check to see if we can do it? why not solve \[.11x+.06(10,000-x)=700\]

OpenStudy (anonymous):

\[\frac{ x*1*6 }{ 100 }+\frac{ \left( 10000-x \right) *1*11}{ 100 }=700\] find x

OpenStudy (misty1212):

or better yet \[11x+6(10,000-x)=70000\]

Nnesha (nnesha):

 that;s how my tec did but i have no idea how he got that equation